OpenStudy (paystub-2):

5p+2=4p-1 solve for p and gve me the steps please:]

OpenStudy (anonymous):

5p+2=4p-1 -4p -2 -4p -2 p=-3

OpenStudy (paystub-2):

how do you check if its right?

OpenStudy (anonymous):

plug in p 5(-3)+2=4(-3)-1 -15+2=-12-1 -13=-13

OpenStudy (paystub-2):

thanks so much can you help me on anymore?

OpenStudy (anonymous):

sure

OpenStudy (paystub-2):

7a+10=2a

OpenStudy (anonymous):

well if i work them all out completely for you it won't help you much. First step seperate the variable. In this case it is "a". You want to have "a" on one side and some number on the other. Try and and tell me what you got

OpenStudy (paystub-2):

I dont need the straight answer right away i just need help getting to the answer cuz when i check it it comes out wrong. the problem is the variable which i cant figure out. Can you set up the problem for me or something?

OpenStudy (anonymous):

Like i said earlier. The variable in your problem is "a". Usually the variable is expressed by some sort of symbol or letter. In this problem you start with: 7a+10=2a Notice the variable "a" is on both sides of the equation and we need it only on one. (either the left or right side it does not matter. What would you do in order to get A on one side of the equation?

OpenStudy (paystub-2):

use the distributive property?

OpenStudy (anonymous):

In this problem the using the distributive property is not needed. try moving 7a to the right side.. what would your equation look like then?

OpenStudy (paystub-2):

10+9a?

OpenStudy (paystub-2):

add 7a to the 2a to get 9a and then add 10 to it.

OpenStudy (anonymous):

leave the 10 be for the moment. In order to move the 7a to the other side you must subtract it from the left and the right 7a+10=2a -7a -7a ( I am not trying to drag this out or anything just figure its better if you actually understand what needs to be done)

OpenStudy (paystub-2):

you must think im done but thats why im the regular class 4 this stuff.

OpenStudy (paystub-2):

meant 2 say dum

OpenStudy (anonymous):

lol naw, just doesnt click for some people... Keep practicing and you will be enlightened eventually :)

OpenStudy (paystub-2):

dats exactly wat i did but i checked my problem it was rong

OpenStudy (anonymous):

do you mind showing me your work so that I can put out where the mistake is?

OpenStudy (anonymous):

point out*

OpenStudy (paystub-2):

10/-5=-5/-5

OpenStudy (paystub-2):

hello....

OpenStudy (anonymous):

sorry lol so you got -2=a

OpenStudy (paystub-2):

oh now i c my mistake it helps when you hav someone like you to be around thanks.

OpenStudy (paystub-2):

its a bit late so i have to go but can we get in contact through this so that you can tutor me again?

OpenStudy (anonymous):

If I am on. I am usually not on much but their are plenty of other people that will gladly help

OpenStudy (anonymous):

there

OpenStudy (paystub-2):

how bout tomorrow morning its just since your just made a coonection w/ u and rly this is my fisrt time trying out this program cuz i found it on the internet. If i do find somebody else hopefully itl be as helpful as you were.

OpenStudy (anonymous):

I have classes all day tomorrow. Just post your question and people will come :P

OpenStudy (paystub-2):

k but ill search 4 u though. bye

OpenStudy (paystub-2):

;)